OpenStudy (anonymous):

Algebra Help Pleaseee 3x+2y=7

OpenStudy (cwrw238):

what is the question?

OpenStudy (anonymous):

"Solve for x" 3x+2y=7 @cwrw238

OpenStudy (cwrw238):

cj49 has solved for y

OpenStudy (cj49):

y=7-3x/2 when x=1 y is 2 i.e 7-3(1)/2 =7-3/2 =4/2 =2

OpenStudy (anonymous):

so x = 2? @cj49

OpenStudy (cwrw238):

to solve for x 3x + 2y = 7 subtract 2y from both sides 3x = 7 - 2y now divide both sides by 3 x = 7 - 2y ----- 3

OpenStudy (cj49):

no x =1 and y=2

OpenStudy (anonymous):

ohhhhh i understand okay thanks guys (:

OpenStudy (cwrw238):

to find the value of y for given value of x plug x into the formula

OpenStudy (cwrw238):

x = 2 y = 7 - 2*2 ----- = (7-4) / 3 = 3/3 = 1 3

OpenStudy (cj49):

np..

OpenStudy (cwrw238):

oops - sorry i got that wrong x = 7 - 2y ----- 3 when y = 2 x = 7 - 2*2 ------ = 3/3 = 1 3

OpenStudy (cwrw238):

I mixed up my y's and x's !